only 20 percent of the towels were wet. if 520 towels were dry,how many towels were there in all?
Answer by checkley77(12569)

.20X=X-520
.20X-X=-520
-.8X=-520
X=-520/.8
X=650 TOTAL NUMBER OF TOWELS.
pROOF:
.20*650=650-520
130=130